In all fours, 6 cards are dealt from a 52-card deck to each of 4 players and then

a card is dealt face up to determine the trump suit. The four players form two teams of two

players each. Suppose that a 10 of clubs is dealt as the face up card. Suppose also, that the

first player was dealt with a Jack and 3 cards with ranks less than 10, of the trump suit.

What is the probability that the second player dealt gets the highest trump in the deck i.e.

Ace of clubs?


1
Expert's answer
2021-09-21T09:44:31-0400

There are a total of 52 cards, the first player got 6 cards, 10 of clubs is dealt as the face up card, out of it so 45 cards remain (52- 6-1= 45)

Clubs is the trump suite right now and the first player got 4 cards out of it so 8 of clubs remain.

So the probability to get a trump is 8/45= 0.1777

The probability to Ace of clubs = 1/8*0.1777=0.125*0.1777=0.0222125

Of 6 cards are dealt from deck to each players.

So the probability that the second player dealt gets the highest trump in the deck = 6*0.0222125= 0.133275